# Routing and wavelength assignment in quantum key distribution networks: power control heuristics for quantum-classical multiplexing

### v1

@@ V-A Set up
We also normalize the launch power of the classical channels with respect to the maximum launch power. This maximum power launch is determined by ensuring that the longest possible path for that particular connection achieves a minimum SNR of $20$ dB for a normalized noise power $N=1$. Once the RWA algorithms identify the candidate path, they compute the required transmission power for that particular path and normalize it to the maximum launch power. Consequently, both $P_{tx,q}$ and $P_{tx,c}$ are normalized at their respective scales. Since the transmission power of quantum and classical lightpaths differ in magnitude, this difference is accounted for in the $N_{c}$ term.
@@
We use the request blocking ratio and the average SNR as the figures of merit to evaluate the performance of the algorithms. The total number of requests arriving to the network varies from $10$ to $100$ with an increment of $10$ requests. For each topology, $500$ simulations are run. Figures include $95\%$ confidence intervals. The averages in the experiments are fairly stable and the confidence intervals are difficult to appreciate in some of the figures.
V-B Results
Figure 3: Difference in the blocking ratio obtained by the RWA algorithms when the Power Control protocol is applied, considering the actual shared fiber length $L_{i}$.
Figure 4: Difference in the blocking ratio obtained by the RWA algorithms when the Power Control protocol is applied, considering the shared effective length $L_{\text{eff}}$.
Figure 5: Difference in the SNR achieved by the RWA algorithms when the Power Control protocol is applied, considering the shared effective length.
Figures 3 and 4 show the blocking ratio of the algorithms with transmission power control compared to the same RWA algorithms that do not apply this control protocol, using $L_{i}$ and $L_{\text{eff}}$ respectively. It can be observed that the results are very similar for both shared distance calculations. Therefore, we will analyze the results obtained using $L_{\text{eff}}$.
Among the algorithms with no Transmission Power Control protocol, KSP-FF demonstrates the best performance in terms of accepted requests, achieving a $0\%$ rejection rate for total requests under $50$ and outperforming the quantum-oriented MQDO and MQCCO by up to $2$ orders of magnitude. QTD exhibits the highest blocking ratio, with blocking ratios above $10^{-1}$, as it consumes the most resources. When the Power Transmission Protocol is applied, all the algorithms except QTD show improved performances. This is expected. In QTD no fiber link shares quantum and classical channels. KSP-FF reduces the blocking ratio by up to $2$ orders of magnitude, while MQDO and MQCCO can improve their performance by up to $1$ order of magnitude. Therefore, effective control of the power transmission of the classical channels enhances the performance of the QKD network.
@@ V-B Results
Figure 5 shows the achieved QSNR of the established quantum channels. The QSNR is consistently above the $15$ dB threshold across all algorithms. Without active transmission power control, the QSNR decreases, as expected, due to increased noise on the quantum lightpaths caused by the higher transmission power of the classical channels. A smaller QSNR is related to a higher QBER which, in turn, means the QKD protocol has a slower key generation rate. Noisier channels produce secret key bits more slowly.
Specifically, KSP-FF achieves a QSNR between $24-27$ dB without power control and improves by up to $3$ dB for the higher number of requests when power control is active. MQDO and MQCCO exhibit similar behavior, with improvements up to $2$ dB. QTD behaves the same with and without power control (there are no shared paths).
When power control is not active, KSP-FF yields lower quality channels compared to quantum-oriented RWA algorithms. While quantum-oriented algorithms can enhance the quality of quantum channels by up to $2$ dB, they do so at the cost of rejecting more requests than KSP-FF. However, with active power control, the quality of channels established by the algorithms is nearly identical.
@@
These experiments suggest proper control of classical transmission power can enhance signal quality and enables the use of classical RWA algorithms for solving the resource allocation problem.
V-C Performance of mixed traffic scenarios.
We also investigate the power control protocol over RWA algorithms in mixed traffic scenarios, i.e., networks in which both quantum and classical requests are present. We consider a network trying to establish QKD links over a network with classical traffic outside our control. Specifically, we analyze the performance of the network for $90$ total requests. We define the classical traffic load as the fraction of the total requests that belong to purely classical traffic. This classical load varies from 00 (representing all the request are quantum) to $1$ (representing all the requests are classical). All other simulation parameters remain unchanged and results are represented with $95\%$ confidence intervals.
Figure 6 shows the blocking ratio of the RWA algorithms with power control protocol and when this protocol is not active. As in 4, when the classical traffic load is 00, the best performing algorithm is KSP-FF with power control protocol. However, when the classical traffic load increases, the performance of KSP-FF with and without power control deteriorates. This decline is due to the potential interference introduced by the increasing number of classical lightpaths, which degrades the QSNR of the established quantum channels. The algorithms are designed to cause a rejection of the classical connection request if they degrade the quality of the established quantum channels below a certain threshold, hence the increase on the blocking ratio.
On the other hand, MQDO and MQCCO improve their performance because they are designed to minimize the shared distance with established quantum channels, thereby reducing interference and degradation of these lightpahts’ quality. Considering that each quantum connection requires the establishment of three additional classical lightpaths, an increase in the classical load implies a reduction in the total number of quantum requests and the associated classical connections, leading to a decrease in the network resource consumption and, consequently, a reduction of the blocking ratio compared to KSP-FF by almost $2$ orders of magnitude. Moreover, applying the power control protocol, we can reduce the blocking ratio by an order of magnitude compared to the same algorithms without power control. Finally, the worst-performing algorithm is QTD, although for the higher number of classical requests, finding a disjoint path for the classical channels can improve the blocking ratio with respect to KSP-FF.
The QSNR achieved by these algorithms is shown in figure 7. The power control protocol enhances the performance of KSP-FF, MQDO, and MQCCO by up to $2$ dB, and QTD by almost $1$ dB (fot power control restricted to the classical channels in the quantum requests). In QKD networks with classical traffic, KSP-FF with power control can slightly improve the quality of the established channels, although at the expense of increasing the blocking ratio. Also, MQCCO and MQDO achieve almost identical QSNRs for the established quantum channels and equal blocking ratio, meaning that the extra effort in accounting for all the established classical channels that MQCCO makes when selecting candidate paths does not provide benefits in terms of either blocking ratio or quality of the quantum channel.
Figure 6: Difference in the blocking ratio obtained by the RWA algorithms when the Power Control protocol is applied, considering the actual shared fiber length $L_{i}$.
Figure 7: Difference in the blocking ratio obtained by the RWA algorithms when the Power Control protocol is applied, considering the shared effective length $L_{\text{eff}}$.
VI Conclusions
@@
In this study, we evaluate the performance of a power control algorithm over a traditional RWA algorithm like KSP-FF and other quantum-oriented RWA algorithms proposed in the literature. This protocol aims at reducing the launch power of the classical lightpaths in order to reduce the noise introduced to the quantum connections. Our study accounts for the undesired noise that classical channels potentially introduce into the quantum channels, proposing a QSNR formula inspired by the SiNR in wireless networks. The study demonstrates that implementing a power control protocol effectively mitigates the noise introduced to quantum connections. This approach enhances the QSNR by reducing interference, thereby improving the quality of quantum channels.
@@
Lastly, in mixed traffic scenarios, where both quantum and classical connection requests are present, the study finds that the increasing number of established classical lightpaths can degrade the quality of the quantum channels despite the use of power control. Quantum-oriented RWA algorithms, which consider the unique requirements of quantum communication, are necessary to effectively manage network resources and maintain high QSNR. These algorithms are crucial for minimizing the impact of classical channels on quantum communication, ensuring the reliability and security of the network.
VII Acknowledgments
Lidia Ruiz has been funded by the European Union NextGeneration UE/MICIU/Plan de Recuperación, Transformación y Resiliencia/Junta de Castilla y León. J.C. Garcia-Escartin has been funded by the European Union NextGeneration UE/MICIU/Plan de Recuperación, Transformación y Resiliencia/Junta de Castilla y León and the Spanish Ministerio de Ciencia e Innovación (MCIN), project PID2020-119418GB-I00.

### v2

@@ 6.1 Set up
We also normalize the launch power of the classical channels with respect to the maximum launch power. This maximum power launch is determined by ensuring that the longest possible path for that particular connection achieves a minimum power of $36.8\text{\,}\mathrm{dBm}$ at the receiver. For realistic detectors and noise levels, this power has shown to be enough for $1.25\text{\,}\mathrm{G}\mathrm{bit}\mathrm{/}\mathrm{s}$ data rates in a fiber network without amplifiers [36]. For power control, the RWA algorithms identify the candidate path and they compute the required transmission power for that particular path and normalize it to the maximum launch power. Consequently, both $P_{tx,q}$ and $P_{tx,c}$ are normalized at their respective scales. Since the transmission power of quantum and classical lightpaths differ in magnitude, this difference is accounted for in the $\gamma_{nl}$ term and in the calculations giving the QSRN and the key rate.
Lastly, the detector quantum efficiency is set to $\eta=0.1$, and the visibility of the interferometer at the receiver is assumed to be $V=0.98$. For these parameters, and considering a BB84 protocol, we set $P_{tx,q}=\eta t$, which takes into account the effects of imperfect detection and assumes we have a mean photon number at the transmitter equal to the transmissivity of the quantum channel $t<1$, chosen to optimize the final key rate [35].
@@
We use the request blocking ratio and the average SNR as the figures of merit to evaluate the performance of the algorithms. The total number of requests arriving to the network varies from $10$ to $100$ with an increment of $10$ requests. For each topology, $50$ simulations are run. Figures include $95\%$ confidence intervals. The averages for different runs of the experiments are fairly stable and the confidence intervals are difficult to appreciate in some of the Figures for the relevant scale. In some case, variability is reduced with a higher network traffic (see, for instance, Figure 2).
6.2 Results
Figure 2: Blocking ratios obtained by the RWA algorithms with and without power control.
Figure 3: QSNR of the established QKD channels using the proposed algorithms (left) and average Secret Key Rate in bits per pulse (right).
Figure 2 shows the blocking ratio of the algorithms with transmission power control in Gabriel topologies with 10 nodes. The blocking ratio is reduced by up to an order of magnitude for the RWA algorithms that implement the power control protocol. This result demonstrates that reducing the power control of the classical channels helps to meet the QSNR requirements, thereby decreasing the blocking ratio and enhancing the performance of the QKD network. This improvement is not observed for the QTD algorithms, with blocking ratios above $10^{-1}$, since they are designed to find separate routes for QKD and classical channels. Therefore, the blocking ratio is primarily due to lack of available resources for establishing new channels, rather than failing to meet the QSNR requirements as in the other algorithms.
@@ 6.2 Results
Figure 3 shows the achieved QSNR (left) and the average Secret Key Rate, SKR, (right) for the quantum channels that are established. The average QSNR is consistently above $15\text{\,}\mathrm{dB}$ across all algorithms. Without active transmission power control, the QSNR decreases, as expected, due to increased noise on the quantum lightpaths caused by the higher transmission power of the classical channels. A smaller QSNR is related to a higher QBER which, in turn, means the QKD protocol has a slower key generation rate.
In our results, all the Figures show the final key bit rate per pulse following the estimation in Equation (5), which can be related to the total secret key rate in bits per second [43]. We have considered a system with detector efficiency $\eta=0.1$ and attenuated coherent pulses with an average photon number $\mu=t$, where $t$ is the transmissivity of the quantum channel and has been chosen to optimize the final key rate [35]. In that case, the expected number of photon detections per pulse, considering a maximum distance of $60\text{\,}\mathrm{km}$ and our attenuation constant in the O band, is $n=\eta t^{2}=\eta 10^{-0.2\alpha_{qdB}}\approx 1.45\cdot 10^{-4}$. Typical QKD lasers can reach pulse repetition rates around $10\text{\,}\mathrm{GHz}$, which would give a photon detection rate of, at least, $1.45\text{\,}\mathrm{MHz}$. If we consider this is the case and that the limiting factor for protocol detections is the detector dead time, which we can set in the microsecond range, the detection rate can be around $1\text{\,}\mathrm{MHz}$. For these settings, the final secret key rate in our simulations would be in the range of a few kilobits per second.
Regarding the performance of each heuristic, KSP-FF achieves a QSNR between 16-$19\text{\,}\mathrm{dB}$ without power control, with an improvement of up to $5\text{\,}\mathrm{dB}$ observed for higher number of requests when power control is enabled. Both MQDO and MQCCO exhibit similar trends; MQDO without power control delivers a QSNR advantage of $1\text{\,}\mathrm{dB}$ over MQCCO and $2\text{\,}\mathrm{dB}$ over SP-FF at low request volumnes. For both MQCCO and MQDO, the QSNR improves up to $5$ dB with power control.
When power control is inactive, all of the algorithms yield lower quality channels compared to quantum-oriented RWA algorithms. The lower quality of the quantum channels translates into slower key generation rate as shown in Figure 3 (right). The algorithms without power control achieve lower average key generation rates compared to the algorithms with power control. In this case, the QTD algorithm achieves the fastest generation rate since the interference caused by the classical channels in the QKD channels is the lowest among the proposed algorithms, but at the cost of a higher blocking probability.
Figure 4: Percentage of power consumption reduction for the RWA algorithms with enabled power control
Lastly, enabling a power control protocol can save up to $80\%$ of energy for all algorithms, as shown in Figure 4. Power consumption is calculated from the mean launch power of the established channels. Subsequently, the power savings are calculated using the following formula:
$\frac{P_{tx,noPC}-P_{tx,PC}}{P_{tx,noPC}}.$ (9)
This expression quantifies the reduction in the average launch power consumption of the established channels when the power control feature is active $P_{tx,PC}$, relative to the consumption when the feature is inactive, $P_{tx,noPC}$. The greatest savings happen for shortest path firtst fit with power control. This is logical, as shorter paths mean there are less losses and a larger excess power with respect to the threshold at the detector. The quantum-oriented heuristics choosing longer paths to avoid contamination into the quantum signals have a slightly lower power reduction.
@@
These experiments suggest proper control of classical transmission power can enhance signal quality and enables the use of classical RWA algorithms for solving the resource allocation problem, while reducing operational expenditures due to the reduction of power consumption.
Figure 5: Blocking ratio and SKR for random Gabriel topologies of 2020 nodes.
Results present a similar trend when the number of nodes is increased to $20$. The blocking ratio is lower, as shown in Figure 5 (right). Since the network topologies are larger while the traffic volume remains unchanged, the blocking ratio decreases due to the increased availability of resources and the greater number of possibilities for accommodating classical channels without interfering with the QKD channels. On the other hand, RWA algorithms with power control achieve a better blocking ratio, however, the improvement is less pronounced compared to the $10$-node topologies, likely because the blocking ratios are already significantly lower. Having a larger network allows for better alternative routes and less connection requests are rejected. In this scenario with a larger network, we also see a higher variance and the confidence intervals in the simulations are quite visible in the blocking probability, particularly for a smaller number of requests, when the particular topology of the random network is more important.
The SKR also improves with respect to the $10$-node topologies, obtaining key rates of up to $9\cdot 10^{-3}$ bits/pulse for SP-FF, MQDO and MQCCO and higher than $1\cdot 10^{-2}$ for QTD. Also, results show that when the traffic volume increases, RWA algorithms without power control produce slower key rates. Hence, QKD networks must have low traffic requirements in order to maintain high-quality QKD connections.
Lastly, we present the results for $20$-node random Waxman topologies. The chosen $\alpha$ and $\beta$ parameters of the Waxman distribution have been selected to create networks with a medium density of connections. In this case, the blocking probability increases with respect to the blocking probability in the $20$-node Gabriel topologies, as shown in Figure 6 (left). This result suggests that the generated Waxman networks have more difficulties for accomodating the channels and meeting the QSNR requirements. For this reason the key generation rate is also lower compared to the Gabriel graphs, as shown in Figure 6 (right), where the key rates are slightly higher than $2\cdot 10^{-3}$ bits/pulse. Interestingly, enabling power control does not result in a significant improvement in the key generation rate, which points to the limited availability of paths that meet the QSNR requirements as the key factor for the blocking ratio and the lack of quality in the QKD connections. However, power control does reduce the blocking ratio, up to an order of magnitude under high load conditions.
Figure 6: Blocking ratio and SKR for random Waxman topologies of 2020 nodes.
6.3 Performance of mixed traffic scenarios.
We also investigate power control in RWA algorithms in mixed traffic scenarios, i.e., networks in which both quantum and classical requests are present. We consider a network trying to establish QKD links over a network with classical traffic outside our control. In order to do that, we study the effects of an increasing number of classical traffic demands. With these simulations, we try to understand how to balance network resources between classical and quantum channels.
Specifically, we analyze the performance of the network for $90$ total requests, which gives a flexible scenario with enough classical channels to be interesting, but not too many to prevent efficient QKD. We define the classical traffic load as the fraction of the total requests that belong to purely classical traffic. We consider these classical channels are confined to our network and that the signals have never crossed an amplifier. Additionally, we suppose these additional channels are not under our control and they do not follow the rules of our heuristics. Their route is fixed, but we can attenuate them (add power control).
The classical load varies from 0 (representing all the request are quantum, including the request for classical service channels) to $1$ (representing all the requests are classical). All other simulation parameters remain unchanged and results are represented with $95\%$ confidence intervals.
Figure 7: Blocking ratio in hybrid networks for $20$-node random Gabriel topologies.
Figure 8: QSNR and SKR hybrid networks for $20$-node random Gabriel topologies.
Figure 7 shows the blocking ratio of the RWA algorithms with power control protocol and when this protocol is not active for $20$-node random Gabriel topologies. We only show the $20$-node results to avoid repetition, since the trends are similar for the $10$-node random Gabriel graphs and the Waxman graphs.
As in Figure 2, introducing power control results in a lower blocking ratio, although the effect is not so significant, at least for the selected traffic load, as shown in Figure 7. Additionally, the blocking ratio also decreases when the classical traffic load increases. Since a classical connection only requires the establishment of one lightpath, the accommodation is easier compared to a quantum request, which requires the establishment of four different lightpaths. We can also see that, for a high fraction of classical traffic, the blocking ratio can be sensitive to the particular quantum connections required and the network topology, as shown by the increased variability evinced by the wider confidence intervals.
In terms of QSNR, Figure 8 (left) confirms that enabling power control significantly enhances the quality of QKD connections, with improvements of up to $6\text{\,}\mathrm{dB}$. This translates into faster secret key generation, as illustrated in Figure 8 (right). In both cases, the best performance—characterized by the highest QSNR and the fastest key generation rate—is achieved by QTD with power control enabled. For the previous scenarios, QTD gave the same results with and without power control. The classical and quantum paths never shared the same fiber. Now, the classical traffic outside our direct control, which is not associated to QKD channels, can be routed on top of an existing quantum link. Adjusting the power of these channels reduces their noise.
Finally, let us analyze the performance of Waxman topologies in hybrid quantum-classical networks. In this case, the blocking ratio is higher than that observed in Gabriel topologies, likely due to the structural properties of the Waxman graphs, which make it challenging to accommodate both quantum and classical channels while meeting the QSNR requirements. Interestingly, enabling the power control feature yields an improvement of up to one order of magnitude in blocking ratio—greater than the improvement observed with Gabriel topologies, as shown in Figure 9. Additionally, MQCCO and MQDO exhibit slightly better performance compared to the other algorithms in this scenario, which is a behaviour that was not observable in Gabriel topologies. These results suggest that certain network topologies may benefit from RWA algorithms tailored to their specific structural characteristics and resource constraints. Furthermore, increasing classical traffic initially raises the blocking ratio; however, once the load is high enough, the number of required quantum lightpaths decreases, facilitating more efficient resource utilization while meeting the QSNR requirements.
Overall, these findings highlight the influence of network topology on quantum-classical coexistence and demonstrate the importance of adaptive routing and power control strategies in optimizing performance.
In terms of QSNR, Figure 10 (right) shows that SP-FF establishes QKD connections with the lowest signal quality. In contrast, MQDO, QTD, and MQCCO achieve QSNR values up to $2\text{\,}\mathrm{dB}$ higher when power control is disabled. Enabling power control further enhances the quality of QKD signals by up to $4\text{\,}\mathrm{dB}$ compared to the inactive case. In this scenario, the differences in QSNR among the algorithms become less significant when power control is active. Regarding the key generation rate, enabling power control results in a slight improvement in the generation speed.
Overall, these results confirm that, while power control has a positive effect on both QSNR and key generation rate, its impact tends to reduce algorithmic differences, making it a key enabler for more consistent QKD performance across routing strategies.
Figure 9: Blocking ratio in hybrid networks for $20$-node random Waxman topologies.
Figure 10: QSNR and SKR hybrid networks for $20$-node random Waxman topologies.
7 Conclusions
@@
In this study, we evaluate the performance of a power control algorithm over a traditional RWA algorithm like KSP-FF and other quantum-oriented RWA algorithms proposed in the literature. This protocol aims at reducing the launch power of the classical lightpaths in order to reduce the noise introduced to the quantum connections. Our study accounts for the undesired noise that classical channels potentially introduce into the quantum channels, proposing a QSNR formula inspired by the signal-to-interference-plus-noise ratio, SiNR, used in wireless networks. The study demonstrates that implementing a power control protocol effectively mitigates the noise introduced to quantum connections. This approach enhances the QSNR by reducing interference, thereby improving the quality of quantum channels.
@@
In hybrid traffic scenarios—where both quantum and classical connection requests coexist—we observed that network topology plays a critical role in performance. The studied graph models exhibited different sensitivities to interference and path availability, suggesting that topology-aware RWA strategies are essential. Moreover, as classical traffic increases, the cumulative effect of classical lightpaths can still degrade quantum channel quality, even with power control in place. In these cases, quantum-aware RWA algorithms such as MQDO, QTD, and MQCCO outperform general-purpose approaches by better accommodating the specific requirements of QKD traffic. Quantum-oriented RWA algorithms, which consider the unique requirements of quantum communication, are necessary to effectively manage network resources and maintain a high key rate. The proposed heuristics for routing and channel allocation minimize the impact of classical channels on quantum communication, ensuring the reliability and security of the network. However, to fully exploit their benefits, they should be assessed for dynamic traffic assignment and in combination with topology-aware and quantum-specific RWA strategies, which are left for future studies.
Overall, our findings confirm that power control is a vital mechanism for enhancing QKD performance in mixed quantum-classical networks. The proposed routing and allocation heuristics contribute to minimizing the impact of classical interference, ensuring reliable and secure quantum communication even in complex, high-load network environments.
Acknowledgments
Lidia Ruiz has been funded by the European Union NextGeneration UE/MICIU/Plan de Recuperación, Transformación y Resiliencia/Junta de Castilla y León n and Junta de Castilla y León (Consejería de Educación/FEDER) Project VA184P24. J.C. Garcia-Escartin has been funded by the European Union NextGeneration UE/MICIU/Plan de Recuperación, Transformación y Resiliencia/Junta de Castilla y León and Junta de Castilla y León (Consejería de Educación/FEDER) Project VA184P24. Financial support of the Department of Education, Junta de Castilla y León, and FEDER Funds is gratefully acknowledged (Reference: CLU-2023-1-05).